A wildlife researcher is tracking a flock of geese. The geese fly 5.0 km due west, then turn toward the north by 45 ∘ and fly another 3.5 km .
Answer #1

(a) The x-component of total distance

Sx = 5 km + 3.5*cos 45 degree km = 7.47 km

(b) Displacement, S = sqrt(Sx^2 + Sy^2)

Sy = 3.5*sin 45 degree = 2.47 km

S = sqrt((7.47)^2 + (2.47)^2))

= 7.87 km